1.One baked potato provides an average of 0.0310 g of vitamin C and a 5.0 lb bag of potatoes contains 15 potatoes. If you consume 2.4 ounces of potatoes, how many miligrams of vitamin C did you consume?
2.One serving of olive oil contains 15mL. If you use 25.5 mL of olive oil in a recipe, how many pints is that?
3.Convert 5.26 cm3 to gallons.
4.A football player ran 18 yards on a play. How many inches did they run?
1
Expert's answer
2018-04-13T08:53:00-0400
1. 1 baked potato – 0.0310 g of vitamin C

5.0 lb bag of potatoes – 15 potatoes

5.0 lb = 80 ounces

2.4*15/80 = 0.45 potato

Vitamin C = 0.45*0.0310 g = 0.01395 g = 13.95 mg

2. 1 ml = 0.0021134 pint

25.5 ml = 0.0539 pints

3. 5.26 cm3 = 0.00139 gallons

4. 1 yard = 36 inches

18 yards = 648 inches